OpenStudy (anonymous):

Use trig identites to transform the left side of equation into the right. (1+cos theta)(1-cos theta)= sin^2theta

OpenStudy (aravindg):

1+cos theta =2cos^2 theta/2

OpenStudy (anonymous):

multiply on the left, and get \[1-\cos^2(\theta)\] then it should be easy enough

OpenStudy (aravindg):

1-cos theta=2sin^2 theta/2

OpenStudy (anonymous):

is it clear that \[1-\cos^2(\theta)=\sin^2(\theta)\]?

OpenStudy (anonymous):

ooohh. yea. I see.

OpenStudy (anonymous):

kk good

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

yw

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

use identity ---> a^2 - b^2 for lhs then prove using cos2 + sin2 = 1

OpenStudy (anonymous):

(1+cosθ)(1−cosθ)=(1−cosθ+cosθ−cos2θ) = sin2theta